OpenStudy (hap1236):

Use the formula for the sum of the first n terms of a geometric sequence to solve. Find the sum of the first 8 terms of the geometric sequence: -8, -16, -32, -64, -128, . . . .

OpenStudy (hap1236):

@FaiqRaees

jhonyy9 (jhonyy9):

than you assume the first with 3rd how many get ? and so too the second with 4th how many get ?

OpenStudy (hap1236):

i dont know @jhonyy9

OpenStudy (hap1236):

@jhonyy9 help

OpenStudy (hap1236):

@ganeshie8

jhonyy9 (jhonyy9):

-8-32=-40 and -16-64 = -80 so what is your opinion about this ?

OpenStudy (hap1236):

i got -248

OpenStudy (hap1236):

A. -2003 B. -2040 C. -2060 D. -2038

OpenStudy (hap1236):

but these are my choices

jhonyy9 (jhonyy9):

-8-16-32-64-128-256-512-1024 = -8-32 =-40 -16-64=-80 -128-512=-640 -256-1024=-1280 ---------------------- assume these = -2040

jhonyy9 (jhonyy9):

so choice B. is true

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (hap1236):

thankyou.. can you help me with ssome others

jhonyy9 (jhonyy9):

welcome just post on the open questions column and tag me there please